AnishGC:

Write the Formula of SinC + SinD = ? And Sin3theta = ?

KyledaGreat:

\[SinC + Sin D = 2\sin \frac{ C+D }{ 2 } \cos \frac{ C-D }{ 2}\] \[Sin 30 = 3\sin \theta - 4\sin^{3}\theta \]